OpenStudy (anonymous):

Factor the expression. 16m2 – 24mn + 9n2 A. (4m – 3n)(4m + 3n) B. (16m – 3n)(m + 3n) C. (4m – 3n)2 D. (4m + 3n)2

OpenStudy (anonymous):

C is the right answer

OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (paxpolaris):

\[\Large \left( a+b \right)^2= a^2+2ab+b^2\] \[\Large \left( a-b \right)^2= a^2-2ab+b^2\] \[\Large \left( a+b \right)\left( a-b \right)= a^2-b^2\]

OpenStudy (paxpolaris):

come on try it your self...

OpenStudy (anonymous):

(4m+3n)2 (4m)^2-2(4m)(3n)+(3n)^2 16m2-24mn+9n2